# Bias Beyond English: Evaluating Social Bias and Debiasing Methods in a Low-Resource Setting

### v1

@@ 3.2 Dataset Construction
The dataset contains 1,508 examples, covering various bias types and measures model bias by comparing the likelihood of stereotypical vs. non-stereotypical sentences. Previous research shows that many widely used language models favor stereotypical sentences in English, revealing internal biases. However, CrowS-Pairs is based on American sociocultural contexts, which limits its applicability to other languages and cultures. To make bias evaluation more representative, this study selects four primary bias types for in-depth analysis: Gender, Race-Color, Nationality, Religion. After filtering the dataset, we obtained 1,042 sentences, adapted for evaluating bias in non-English languages.
We translated CrowS-Pairs into four languages: Chinese (zho), Russian (rus), Indonesian (ind), and Thai (tha) using the Google Translate API. These languages were chosen based on Common Crawl Crawl (2024), a large-scale web dataset that reflects the availability of online resources in different languages. Table 1 shows the proportion of web content in these languages. According to linguistic resource classifications in NLP (Joshi et al., 2020), Chinese and Russian are considered high-resource, while Thai and Indonesian are low-resource languages, despite having millions of native speakers.
@@ 3.3.1 Setting
In this study we selected six pre-trained models: mBERT, XLM-RoBERTa, XGLM, Gemma 3, Qwen 2.5 and LLaMA 3. These models are widely used particularly in multilingual tasks. For the details of the models, please refer Appendix B. Note that all language models in this list support all evaluated languages, except Qwen 2.5, which has reported to support the other four language but not Thai, as indicated in its technical documentation (Yang et al., 2025). Nevertheless, we still report the bias scores calculated for Qwen in Thai as per our evaluation methodology; however, the reader should be advised to keep the information in mind that qwen 2.5 model lacks proper support for Thai language, as will be discussed in later sections.
@@ 4 Cross-lingual Transfer Debiasing
We have first examined how biases are presented across languages, and now we proceed to investigate whether other languages can benefit from English corpus-based debiasing methods. We use XLM-RoBERTa in this case as an example because it has high bias, and there is a significant gap between English and other languages in terms of both bias severity and representational alignment. This discrepancy makes XLM-RoBERTa an illustrative case for investigating whether debiasing techniques developed for English can transfer effectively to other languages. To quantify the effect of debiasing techniques, we calculated the degree of bias reduction for each debiasing method as the relative percentage reduction:
@@ 4.1 Data
Following the similar experiments done by Reusens et al. (2023a), we select our debiasing experiment data from the Wikipedia dataset Meade et al. (2022), which contains cleaned multilingual full-text Wikipedia articles. The dataset is built based on Wikipedia data dumps, with unnecessary elements such as citations and other non-main content removed from the articles. For the purpose of this research, we selected 10% of the data from Wikipedia’s extensive database for experimentation. Through this approach, we obtained unsupervised data from 514,084 articles as a sufficient sample size for our bias assessment.
@@ 4.2 Methods
We adpoted three key strategies: CDA (Zmigrod et al., 2019; Webster et al., 2020), DO and SenDeb (Liang et al., 2020). Beyond the three, there are also other approaches such as Self-Debias, INLP and DR. However, since Self-Debias is a post-hoc text generation debiasing procedure, it cannot be used as a debiasing technique for downstream natural language understanding tasks; INLP, DR and SenDeb are all projection-based debiasing techniques, therefore we only chose to experiment with SenDeb. In order to measure the effectiveness of the debiasing methods, we conducted experiments using multilingual models and reported the bias decrease percentage before and after the debiasing experiments. For a complete methodological-level description, please refer to appendix C.
@@ 5 Conclusion and Future Directions
Bias in language models has garnered widespread attention in English, but this challenge is especially pronounced for low-resource languages, which often lack sufficient training data. This study focused on two main aspects: Evaluating bias in a multilingual setting and demonstrating the cross-lingual transferability of debiasing methods, offering new perspectives for ensuring fairness in multilingual models.
In this study we proposed a new evaluation metric to enable fair comparisons between different models, constructed a multilingual bias evaluation dataset, consisting of languages with different resource status, and benchmarked major multilingual language models. We have found that for LLMs transfer learning of fairness is much better for non-English language compared to previous models, and suggested that there exists a nuance of social bias when dealing with different linguistic and cultural backgrounds. Our study demonstrates that high-resource language debias methods can be effectively transferred to low-resource languages for bias mitigation. SenDeb emerged as the most effective technique, suggesting that bias subspaces may share cross-linguistic properties, enabling cross-language debiasing. This finding opens new possibilities for developing universal debiasing methods across diverse languages.
@@
  • Enhance model performance in low-resource languages with larger-scale training and fine-tuning approaches.
@@
Our bias evaluation dataset was constructed through translation, primarily based on an American cultural and linguistic context. This could introduce unfair comparison given that the datasets do not necessarily reflect the sociocultural realities of other languages, potentially affecting the accuracy of our cross-lingual bias evaluations. The bias evaluation and debias methods in this study were designed for masked language models. While we approximate their application to causal language models (autoregressive transformers), there may be inherent differences in how bias manifests and propagates in these models. While our study suggests that bias subspaces share cross-linguistic properties, the extent to which this holds across all languages remains an open question. The effectiveness of debiasing methods may vary for languages with fundamentally different linguistic structures or those underrepresented in pretraining corpora. Further research is needed to validate and refine our findings across a wider linguistic and typological spectrum.
Acknowledgment
We would like to thank Kajetan Dymkiewicz, Songbo Hu, Caiqi Zhang, Sanhanat Sivapiromrat and all those who generously contributed their time to proofread this manuscript and provide valuable comments and suggestions.

### v2

@@ 3.2 Dataset Construction
The dataset contains 1,508 examples, covering various bias types and measures model bias by comparing the likelihood of stereotypical vs. non-stereotypical sentences. Previous research shows that many widely used language models favor stereotypical sentences in English, revealing internal biases. However, CrowS-Pairs is based on American sociocultural contexts, which limits its applicability to other languages and cultures. To make bias evaluation more representative, this study selects four primary bias types for in-depth analysis: Gender, Race-Color, Nationality, Religion, which are relatively generalizable across different cultural and linguistic contexts. After filtering the dataset, we obtained 1,042 sentences, adapted for evaluating bias in non-English languages.
We translated CrowS-Pairs into four languages: Chinese (zho), Russian (rus), Indonesian (ind), and Thai (tha) using the Google Translate API. The languages were chosen based on Common Crawl Crawl (2024), a large-scale web dataset that reflects the availability of online resources in different languages. Table 1 shows the proportion of web content in these languages. According to linguistic resource classifications in NLP (Joshi et al., 2020), Chinese and Russian are considered high-resource, while Thai and Indonesian are low-resource languages, despite having millions of native speakers. The translation tool was chosen considerting both the effectiveness and budget-resource constraints. Despite the resulting corpus being artificial and translation-based, we will see in the later section that the experimental results still reveal meaningful patterns, demonstrating the value of the evaluation albeit limitations.
@@ 3.3.1 Setting
In this study we selected six pre-trained models: mBERT, XLM-RoBERTa, XGLM, Gemma 3, Qwen 2.5 and LLaMA 3. These models are widely used particularly in multilingual tasks. For the details of the models, refer Appendix B. Note that all language models in this list support all evaluated languages, except Qwen 2.5, which has reported to support the other four language but not Thai, as indicated in its technical documentation (Yang et al., 2025). We still report the bias scores calculated for Qwen 2.5 in Thai as per our evaluation methodology; however, the reader should be advised to keep the information in mind that Qwen 2.5 model lacks proper support for Thai language, as will be discussed in later sections.
@@ 4 Cross-lingual Transfer Debiasing
We have first examined how biases are presented across languages, and now we proceed to investigate whether other languages can benefit from English corpus-based debiasing methods. We use XLM-RoBERTa in this case as an example because there is a significant gap between English and other languages in terms of both bias severity and representational alignment. This discrepancy makes XLM-RoBERTa an illustrative case for investigating whether debiasing techniques developed for English can transfer effectively to other languages. To quantify the effect of debiasing techniques, we calculated the degree of bias reduction for each debiasing method as the relative percentage reduction:
@@ 4.1 Data
Following the similar experiments done by Reusens et al. (2023a), we select our debiasing experiment data from the Wikipedia dataset Meade et al. (2022), which contains cleaned multilingual full-text Wikipedia articles. For the purpose of this research, we selected 10% of the data from Wikipedia’s extensive database for experimentation. Through this approach, we obtained unsupervised data from 514,084 articles as a sufficient sample size for our bias assessment.
@@ 4.2 Methods
We adpoted three key strategies: CDA (Zmigrod et al., 2019; Webster et al., 2020), DO and SenDeb (Liang et al., 2020). Beyond the three, there are also other approaches such as Self-Debias, INLP and DR. However, Self-Debias is a post-hoc text generation debiasing procedure and cannot be used as a debiasing technique for downstream natural language understanding tasks; INLP, DR and SenDeb are all projection-based debiasing techniques, therefore we only chose to experiment with SenDeb. For a complete methodological-level description, please refer to appendix C.
@@ 5 Conclusion and Future Directions
This study focused on two main aspects: Evaluating bias in a multilingual setting and demonstrating the cross-lingual transferability of debiasing methods, offering new perspectives for ensuring fairness in multilingual models.
We proposed a new evaluation metric to enable fair comparisons between different models, constructed a multilingual bias evaluation dataset, consisting of languages with different resource status, and benchmarked major multilingual language models. We have found that for LLMs transfer learning of fairness is much better for non-English language compared to previous models, and suggested that there exists a nuance of social bias when dealing with different linguistic and cultural backgrounds. Our study demonstrates that high-resource language debias methods can be effectively transferred to low-resource languages for bias mitigation. SenDeb emerged as the most effective technique, suggesting that bias subspaces may share cross-linguistic properties, enabling cross-language debiasing. This finding opens new possibilities for developing universal debiasing methods across diverse languages.
@@
Our bias evaluation dataset was constructed through translation, using English—specifically American English—as the cultural and linguistic base. While this enabled efficient multilingual expansion, it risks importing U.S.-centric social norms into other languages, potentially skewing bias evaluations and leading to unfair cross-lingual comparisons. Also, in this work we rely on Google Translate to construct new corpus into due to budget and availability constraints, which restricted our ability to ensure culturally nuanced translations. To address the issue mentioned above, future work should move beyond translation-based methods and invest in building culturally grounded datasets developed natively in each language. Collaborating with native speakers and cultural experts will be key to capturing locally relevant social dynamics and ensuring more accurate, equitable bias evaluations.
The bias evaluation and debias methods in this study were designed for masked language models. While we approximate their application to causal language models (autoregressive transformers), there may be inherent differences in how bias manifests and propagates in these models.
While our study suggests that bias subspaces share cross-linguistic properties, the extent to which this holds across all languages remains an open question. The effectiveness of debiasing methods may vary for languages with fundamentally different linguistic structures or those underrepresented in pretraining corpora. Further research is needed to validate and refine our findings across a wider linguistic and typological spectrum.